OpenStudy (anonymous):

Find the perimeter of an equilateral triangle with side length of 12 inches.

OpenStudy (saifoo.khan):

Perimeter is the total boundary. A triangle has 3 sides, 12 + 12 + 12 = answer

OpenStudy (theviper):

\[Area of a Triangle=\LARGE{\color{Red}{\frac{1}{2}}\times base \times height}\]